OpenStudy (anonymous):

Arithmetic Sequence Help Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95. 203 197 168 163

OpenStudy (anonymous):

help please :(

OpenStudy (the_fizicx99):

First find the difference between a1 and a18 95 - (7) / 18-1 102 / 17 = 6 The difference is 6 Use the arithmetic formula to find the 35th term an = a1 + d(n – 1) a_35 = -7 + 6(35-1) a_35 = -7 + 204 a_35 = 197

OpenStudy (the_fizicx99):

Understand?

OpenStudy (anonymous):

thank you @tHe_FiZiCx99

OpenStudy (the_fizicx99):

yw